Janey is twice as old as Peter. The sum of their ages is 5 times Peter's age less 48. How old are Peter and Janey? Can you help me with writing an equation to solve the equation and work out their ages please.

Answer by Maths68(1474)

Let
Present age of Janey = j
Present age of Peter = p


Janey is twice as old as Peter.
j=2p.........(1)
The sum of their ages is 5 times Peter's age less 48.
j+p=5p-48
j=5p-48-p
j=4p-48.........(2)
Put the value of j from (1) to (2)
2p=4p-48
2p-4p=-48
-2p=-48
-2p/-2=-48/-2
p=24
Put the value of p in (1)
j=2p
j=2*24
j=48
Present age of Janey = j = 48 years old
Present age of Peter = p = 24 years old